# StatLLaMA: A multi-stage training framework for building a domain-optimized statistical language model

### v1

@@ Abstract
This study investigates how to efficiently build a domain-specialized large language model (LLM) for statistics using the lightweight LLaMA-3.2-3B family as the foundation model (FM). We systematically compare three multi-stage training pipelines, starting from a base FM with no instruction-following capability, a base FM augmented with post-hoc instruction tuning, and an instruction-tuned FM with strong general reasoning abilities across continual pretraining, supervised fine-tuning (SFT), reinforcement learning from human feedback (RLHF) preference alignment, and downstream task adaptation. Results show that pipelines beginning with a base FM fail to develop meaningful statistical reasoning, even after extensive instruction tuning, SFT, or RLHF alignment. In contrast, starting from LLaMA-3.2-3B-Instruct enables effective domain specialization. A comprehensive evaluation of SFT variants reveals clear trade-offs between domain expertise and general reasoning ability. We further demonstrate that direct preference optimization provides stable and effective RLHF preference alignment. Finally, we show that downstream fine-tuning must be performed with extremely low intensity to avoid catastrophic forgetting in highly optimized models. The final model, StatLLaMA, achieves strong and balanced performance on benchmarks of mathematical reasoning, common-sense reasoning, and statistical expertise, offering a practical blueprint for developing resource-efficient statistical LLMs. The code is available at https://github.com/HuangDLab/StatLLaMA.
Keywords: large language model, domain adaptation, multi-stage training strategy, foundation model, supervised fine-tuning, parameter-efficient fine-tuning, continual pretraining, reinforcement learning from human feedback
@@ 1 Introduction
The development of large language models (LLMs) has profoundly reshaped natural language processing (NLP). Transformer-based architectures, such as BERT [Devlinetal2019bertpre-training], the GPT series [Radfordetal2018improvinglanguage, Radfordetal2019languagemodels, Brownetal2020languagemodels], and the recent LLaMA series [Touvronetal2023llamaopen, Touvronetal2023llama2open, Grattafiorietal2024thellama3], have demonstrated exceptional performance across a wide spectrum of general-purpose NLP tasks. These models have shifted the paradigm from zero-shot and few-shot learning [Brownetal2020languagemodels] toward more sophisticated instruction-following capabilities [Weietal2022finetunedlanguage, Ouyangetal2022traininglanguage]. Built upon large-scale pretraining using massive text corpora, these foundation models [Bommasanietal2022onthe] encapsulate extensive world knowledge and general language understanding.
@@
Supervised fine-tuning (SFT) adapts general-purpose models to specific domains by training on input-output mapping pairs of domain tasks. Traditional implementations of SFT, such as used in BERT, often involve full fine-tuning (FFT), which updates all model parameters. While effective, FFT becomes prohibitively expensive as model sizes grow, making it unsuitable for deploying separate fine-tuned models per downstream task. Consequently, the research community has explored more efficient adaptation strategies. Early approaches like ULMFiT [Howardetal2018universallanguage] introduced concepts such as discriminative fine-tuning and gradual unfreezing, marking the first steps toward parameter-efficient fine-tuning (PEFT). Recent advancements in PEFT, such as adapter tuning [Houlsbyetal2019parameter-efficienttransfer], prefix-tuning [Lietal2021prefix-tuningoptimizing], prompt tuning [Lesteretal2021thepower], and particularly low-rank adaptation (LoRA) [Huetal2021loralow-rank], enable efficient domain adaptation by freezing the majority of model parameters while training only a small number of task-specific parameters. These techniques reduce computational demands and often outperform in-context learning [Liuetal2022few-shotparameter-efficient]. Tools like QLoRA [Dettmersetal2023qloraefficient] further enhance efficiency by integrating quantization strategies. Nevertheless, the key challenge remains: how to inject domain knowledge deeply while preserving general reasoning. This study investigates this central question using PEFT as a foundational component.
@@
Continual pretraining (CoP) [Gururanganetal2020dontstop, Guptaetal2023continualpre-training, Jinetal2022lifelongpretraining] presents another avenue for domain adaptation, allowing the model to continue learning from domain-specific, unlabeled text. While effective for capturing linguistic and structural characteristics of the target domain, our preliminary experiments revealed that CoP alone without strong instruction-following capabilities yields limited improvements in complex reasoning tasks. Instruction tuning [Weietal2022finetunedlanguage, Ouyangetal2022traininglanguage, Chungetal2024scalinginstruction-finetuned, Longpreetal2023theflan, Taorietal2023alpacaa] has emerged as a powerful technique for enhancing LLM usability and task generalization. By training on large collections of instruction–task–response triplets, models learn to better interpret and execute human instructions. From InstructGPT [Ouyangetal2022traininglanguage] to FLAN [Weietal2022finetunedlanguage, Chungetal2024scalinginstruction-finetuned, Longpreetal2023theflan] and Alpaca [Taorietal2023alpacaa], instruction-tuned models, often referred to as “chat models”, have become the standard for interactive AI systems.
@@
Yet, even SFT or CoP cannot fully ensure that a model aligns with human preferences in tasks involving intricate reasoning or subtle judgment. Reinforcement learning from human feedback (RLHF) [Stiennonetal2020learningto, Baietal2022traininga, Azaretal2024ageneral, Fengetal2024towardsanalyzing] addresses this by using human preference rankings to train a reward model [Christianoetal2017deepreinforcement, Ziegleretal2020fine-tuninglanguage], followed by reinforcement learning—often with proximal policy optimization (PPO) [Schulmanetal2017proximalpolicy]—to optimize alignment. While effective, RLHF’s two-stage process is resource-intensive and sensitive to hyperparameter tuning [Zhengetal2023secretsof]. To overcome these limitations, direct preference optimization (DPO) [Rafailovetal2023directpreference] offers a more streamlined approach by optimizing directly on paired preference data using a classification loss function, eliminating the need for explicit reward model training. Another efficient alternative, group relative policy optimization (GRPO) [Shaoetal2024deepseekmathpushing]—a PPO variant, reduces resource requirements by using group-based scoring to estimate policy advantage. This study provides the first empirical evaluation of DPO and GRPO in the statistics domain. High-quality preference data are crucial for both methods. We employed a powerful external LLM (Gemini) as a teacher model to generate response pairs with clear preference distinctions—an approach inspired by knowledge distillation [Hintonetal2015distillingthe, Calderonetal2023asystematic]. Leveraging a stronger model’s judgments provided precise training signals, contributing significantly to our preference optimization success.
@@
Building on these insights, we propose and evaluate a multi-stage training paradigm for developing a domain-specialized statistical LLM. The process begins with the LLaMA-3.2-3B model and incorporates either CoP combined with instruction tuning or SFT using the parameter-efficient LoRA method. These stages are designed to inject core statistical knowledge and task structures into the model. Next, we apply preference alignment using DPO or GRPO to fine-tune the model’s outputs according to human preferences. Finally, we perform carefully calibrated downstream task fine-tuning (DTFT) using high-quality question-answering datasets, aiming to improve domain-specific performance while preserving the model’s core general capabilities. To assess the effectiveness of this approach, we conducted extensive experiments across several benchmark tasks: GSM8K [Cobbe2021trainingverifiers] for mathematical reasoning, AP Statistics multiple-choice questions [APStat2024apstatistics] for statistical knowledge, and ARC [Clarketal2018thinkyou] for commonsense reasoning.
@@ 2.1 Foundation models
Foundation models have become a cornerstone of modern NLP, underpinning many state-of-the-art LLMs. These models—such as BERT [Devlinetal2019bertpre-training], GPT [Radfordetal2018improvinglanguage, Radfordetal2019languagemodels, Brownetal2020languagemodels], and LLaMA [Touvronetal2023llamaopen, Touvronetal2023llama2open, Grattafiorietal2024thellama3]—are built on the Transformer architecture [Vaswanietal2017attentionis] and are pretrained on massive text corpora using self-supervised learning (SSL) [Jingetal2021self-supervisedvisual]. This pretraining stage enables models to learn deep semantic patterns from unlabeled data, developing broad language understanding and general-purpose reasoning capabilities.
@@
This study builds on these insights by using the pretrained LLaMA-3.2-3B model as the foundation for a specialized statistical LLM. The model is built on an optimized Transformer decoder-only architecture, featuring several key improvements over the original Transformer design [Ibe2024unlockinglow-resource]. Whereas the traditional Transformer employs a dual-branch encoder–decoder structure with post-layer normalization, LLaMA retains only the decoder and applies pre-normalization using RMSNorm [Zhangetal2019rootmean] at the input of each sub-layer, enhancing stability in deep model training. For positional encoding, LLaMA replaces traditional additive embeddings with rotary position embeddings [suetal2024roformerenhanced]. The self-attention mechanism in LLaMA employs masked attention and supports both multi-head attention and grouped-query attention [Ainslieetal2023gqatraining], depending on the model scale. LLaMA’s feed-forward network adopts the SwiGLU [Shazeer2020gluvariants] activation function. Collectively, these architectural refinements underpin the high efficiency and performance characteristic of the LLaMA family.
@@ 2.2 Supervised fine-tuning
Supervised fine-tuning (SFT) adapts a pretrained language model to a specific downstream task by learning from labeled data $\mathcal{D}=\{(x,y)\}$, where $x$ is the input and $y$ is the target output. The model parameters $\theta$ are optimized by minimizing the negative log-likelihood (NLL) loss:
@@
While SFT is often implemented as full fine-tuning (FFT), updating all model parameters, this approach becomes prohibitively expensive for large-scale LLMs and is susceptible to catastrophic forgetting [Kirkpatricketal2017overcomingcatastrophic, Lopez-Pazetal2017gradientepisodic, Huangetal2024mitigatingcatastrophic]. Earlier work ULMFiT [Howardetal2018universallanguage] introduced techniques to overcome FFT’s potential risks. ULMFiT’s key strategies included: discriminative fine-tuning (applying different learning rates to different layers, preserving low-level general features), gradual unfreezing (unfreezing layers progressively from top to bottom, allowing adaptation of high-level features before modifying lower-level ones), and slanted triangular learning rates (rapidly increasing the learning rate early in training, then gradually decreasing it to balance adaptation speed with knowledge retention). These methods reduced knowledge degradation and inspired more refined fine-tuning strategies in subsequent large-model training.
@@ 2.3 Parameter-efficient fine-tuning
The core idea of parameter-efficient fine-tuning (PEFT) is to freeze most of the parameters of a pretrained model $\mathbf{\Phi}_{0}$ and introduce only a small set of additional learnable parameters $\mathbf{\Theta}$, typically accounting for less than 1% of the total. During fine-tuning, only $\mathbf{\Theta}$ is updated, while $\mathbf{\Phi}_{0}$ remains unchanged. This approach substantially reduces computational and storage costs and mitigates catastrophic forgetting, as the pretrained parameters are preserved, maintaining the general knowledge acquired during pretraining.
Among various PEFT techniques, low-rank adaptation (LoRA) [Huetal2021loralow-rank] has emerged as one of the most widely used due to its simplicity and effectiveness. LoRA is based on the hypothesis that the weight update $\triangle W$ to a pretrained matrix $W_{0}\in\mathbb{R}^{d\times k}$ can be decomposed into the product of two low-rank matrices:
@@
The success of LoRA has led to numerous extensions. QLoRA [Dettmersetal2023qloraefficient] combines LoRA with 4-bit quantization to reduce memory requirements, enabling fine-tuning of large models on consumer-grade GPUs. DoRA [Liuetal2024doraweight-decomposed] decomposes weights into magnitude and direction, applying low-rank adaptation only to the directional component to potentially improve performance. Further research has examined the impact of LoRA on different layers and modules, such as attention mechanisms and feed-forward networks [Lialinetal2024scalingdown].
@@ 2.4 Continual pretraining
Supervised fine-tuning alone often cannot provide a model with the depth of domain knowledge required to fully capture specialized terminology, linguistic patterns, and background context. To address this limitation, continual pretraining (CoP) has been proposed as an effective strategy for domain knowledge enhancement. Its central idea is to continue training a general-purpose pretrained model on large amounts of domain-specific text—typically unlabeled—using self-supervised learning (SSL) objectives. By extending the objectives proven effective in initial pretraining, CoP allows the model to efficiently absorb knowledge from the new corpus.
Two common approaches dominate CoP. Masked language modeling (MLM), popularized by BERT [Devlinetal2019bertpre-training], randomly masks tokens in the input and trains the model to predict them based on context. This promotes bidirectional contextual understanding, crucial for mastering precise domain-specific terminology. Causal language modeling (CLM)/next token prediction (NTP), used in autoregressive models such as GPT [Radfordetal2018improvinglanguage, Radfordetal2019languagemodels, Brownetal2020languagemodels], predicts the next token given the preceding sequence, enabling the model to learn domain-specific text generation, terminology sequences, and stylistic coherence. Other SSL variants, such as T5’s denoising objective [Raffeletal2020exploringthe] and BART’s text infilling [Lewisetal2019bartdenoising], also provide inspiration for CoP design, though MLM and CLM remain the most widely adopted due to their simplicity and effectiveness.
Recent work has focused on improving CoP’s efficiency and stability. For example, the (re)warming strategy leverages better initialization and learning rate schedules to accelerate adaptation and convergence [Guptaetal2023continualpre-training]. The lifelong pretraining framework enables continuous learning from evolving domain corpus while mitigating knowledge forgetting, addressing the challenge of rapidly changing knowledge bases [Jinetal2022lifelongpretraining].
A key advantage of CoP is that it eliminates the need for costly manual annotation. By immersing the model in large volumes of authentic domain-specific text, CoP enables it to acquire vocabulary, grammar, semantic structures, and even implicit reasoning patterns characteristic of the domain. Empirical evidence shows CoP to be particularly effective in terminology-dense fields, in domains with distinct linguistic styles, or when the target corpus diverges substantially from the original pretraining data [Gururanganetal2020dontstop].
@@ 2.5 Instruction tuning
Instruction tuning [Weietal2022finetunedlanguage, Ouyangetal2022traininglanguage, Chungetal2024scalinginstruction-finetuned, Longpreetal2023theflan, Taorietal2023alpacaa] is a pivotal technique that transforms LLMs from simple “text-completion engines” into versatile AI assistants capable of understanding user intent and following natural language instructions. Its core mechanism is supervised fine-tuning (SFT) on large-scale datasets of instruction–input–output pairs that span diverse task types and formats. Through this process, models learn not only to interpret instructions but also to generalize effectively to new, unseen tasks.
@@
A major factor behind instruction tuning’s success is the scale, diversity, and quality of training data. Early approaches reformatted existing NLP datasets—such as question answering, summarization, and translation—into instruction-based formats, unifying hundreds of task types under a single paradigm [Longpreetal2023theflan]. Later, automated data generation methods further advanced scalability. Self-Instruct [Wangetal2023self-Instructaligning] leveraged stronger teacher models to generate new task instructions and responses from a small set of seed instructions, while XL-Instruct [Iyeretal2025xl-instructsynthetic] extended this approach to build large, high-quality corpora through self-guidance. These methods reduce reliance on manual annotation but introduce challenges such as redundancy, lack of diversity, and stylistic uniformity, which necessitate additional control mechanisms. Other research has incorporated real user–model conversation logs into instruction tuning datasets [Taorietal2023alpacaa]. Although limited in scale and sensitive in terms of privacy and compliance, this type of data reflects authentic usage scenarios and enhances a model’s responsiveness in open-ended dialogue. Its low cost and high application relevance demonstrate the feasibility of instruction tuning under resource-constrained conditions.
In summary, instruction tuning represents a turning point in LLM development, enabling practical human–computer interaction. Its effectiveness derives not only from fine-tuning strategies but also from the diversity and quality of instruction datasets, which may include reformatted NLP tasks, manual annotations, synthetic data generated by teacher models, and real-world interaction data. Together, these sources equip instruction-tuned models with stronger task comprehension and contextual adaptability, laying a robust “base policy” for subsequent alignment techniques such as reinforcement learning from human feedback (RLHF).

### v2

# StatLLaMA: Multi-Stage training for domain-optimized statistical large language models

@@ Abstract
This study investigates how to efficiently build a domain-specialized large language model (LLM) for statistics using the lightweight LLaMA-3.2-3B family as the foundation model (FM). We systematically compare three multi-stage training pipelines—starting from a base FM with no instruction-following capability, a base FM augmented with post-hoc instruction tuning, and an instruction-tuned FM with strong general reasoning abilities—across continual pretraining, supervised fine-tuning (SFT), reinforcement learning from human feedback (RLHF) preference alignment, and downstream task fine-tuning (DTFT). Results show that pipelines beginning with a base FM fail to develop meaningful statistical reasoning, even after extensive instruction tuning, SFT, or RLHF alignment. In contrast, starting from LLaMA-3.2-3B-Instruct enables effective domain specialization. A comprehensive evaluation of SFT variants reveals clear trade-offs between domain expertise and general reasoning ability. We further demonstrate that direct preference optimization provides stable and effective RLHF preference alignment. Finally, we show that DTFT must be performed with extremely low intensity to avoid catastrophic forgetting in highly optimized models. The final model, StatLLaMA, achieves strong and balanced performance on benchmarks of mathematical reasoning, common-sense reasoning, and statistical expertise, offering a practical blueprint for developing resource-efficient statistical LLMs. The code is available at https://github.com/HuangDLab/StatLLaMA.
Keywords: large language model, domain adaptation, multi-stage training strategy, foundation model, supervised fine-tuning, parameter-efficient fine-tuning, continual pretraining, instruction tuning, reinforcement learning from human feedback, downstream task fine-tuning
@@ 1 Introduction
The development of large language models (LLMs) has profoundly reshaped natural language processing (NLP). Transformer-based architectures, such as BERT [15], the GPT series [49, 50, 7], and the recent LLaMA series [63, 64, 17], have demonstrated exceptional performance across a wide spectrum of general-purpose NLP tasks. These models have shifted the paradigm from zero-shot [67] and few-shot [7] learning toward more sophisticated instruction-following capabilities [46]. Built upon large-scale pretraining using massive text corpora, these foundation models [5] encapsulate extensive world knowledge and general language understanding.
@@
Supervised fine-tuning (SFT) adapts general-purpose models to specific domains by training on input-output mapping pairs of domain tasks. Traditional implementations of SFT, such as used in BERT, often involve full fine-tuning (FFT), which updates all model parameters. While effective, FFT becomes prohibitively expensive as model sizes grow, making it unsuitable for deploying separate fine-tuned models per downstream task. Consequently, the research community has explored more efficient adaptation strategies. Early approaches like ULMFiT [24] introduced concepts such as discriminative fine-tuning and gradual unfreezing, marking the first steps toward parameter-efficient fine-tuning (PEFT). Recent advancements in PEFT, such as adapter tuning [23], prefix-tuning [38], prompt tuning [34], and particularly low-rank adaptation (LoRA) [26], enable efficient domain adaptation by freezing the majority of model parameters while training only a small number of task-specific parameters. These techniques reduce computational demands and often outperform in-context learning [41]. Tools like QLoRA [14] further enhance efficiency by integrating quantization strategies. Nevertheless, the key challenge remains: how to inject domain knowledge deeply while preserving general reasoning. This study investigates this central question using PEFT as a foundational component.
@@
Continual pretraining (CoP) [19, 18, 31] presents another avenue for domain adaptation, allowing the model to continue learning from domain-specific, unlabeled text. While effective for capturing linguistic and structural characteristics of the target domain, our preliminary experiments revealed that CoP alone without strong instruction-following capabilities yields limited improvements in complex reasoning tasks. Instruction tuning has emerged as a powerful technique for enhancing LLM usability and task generalization. By training on large collections of instruction-task-response triplets, models learn to better interpret and execute human instructions. From InstructGPT [46] to FLAN [67, 10, 43] and Alpaca [60], instruction-tuned models, often referred to as “chat models”, have become the standard for interactive AI systems.
@@
Yet, even SFT or CoP cannot fully ensure that a model aligns with human preferences in tasks involving intricate reasoning or subtle judgment. Reinforcement learning from human feedback (RLHF) [58, 4, 3, 16] addresses this by using human preference rankings to train a reward model [9, 71], followed by reinforcement learning—often with proximal policy optimization (PPO) [54]—to optimize alignment. While effective, RLHF’s two-stage process is resource-intensive and sensitive to hyperparameter tuning [70]. To overcome these limitations, direct preference optimization (DPO) [51] offers a more streamlined approach by optimizing directly on paired preference data using a classification loss function, eliminating the need for explicit reward model training. Another efficient alternative, group relative policy optimization (GRPO) [55]—a PPO variant, reduces resource requirements by using group-based scoring to estimate policy advantage. This study provides the first empirical evaluation of DPO and GRPO in the statistics domain. High-quality preference data are crucial for both methods. We employed a powerful external LLM (Gemini) as a teacher model to generate response pairs with clear preference distinctions—an approach inspired by knowledge distillation [21, 8]. Leveraging a stronger model’s judgments provided precise training signals, contributing significantly to our preference optimization success.
@@
Building on these insights, we propose and evaluate a multi-stage training paradigm for developing a domain-specialized statistical LLM. The process begins with the LLaMA-3.2-3B model and incorporates either CoP combined with instruction tuning or SFT using the parameter-efficient LoRA method. These stages are designed to inject core statistical knowledge and task structures into the model. Next, we apply preference alignment using DPO or GRPO to fine-tune the model’s outputs according to human preferences. Finally, we perform carefully calibrated downstream task fine-tuning (DTFT) using high-quality question-answering datasets, aiming to improve domain-specific performance while preserving the model’s core general capabilities. To assess the effectiveness of this approach, we conducted extensive experiments across several benchmark tasks: GSM8K [12] for mathematical reasoning, AP Statistics multiple-choice questions [2] for statistical knowledge, and ARC [11] for commonsense reasoning.
@@ 2.1 Foundation models
Foundation models have become a cornerstone of modern NLP, underpinning many state-of-the-art LLMs. These models—such as BERT [15], GPT [49, 50, 7], and LLaMA [63, 64, 17]—are built on the Transformer architecture [65] and are pretrained on massive text corpora using self-supervised learning (SSL) [32]. This pretraining stage enables models to learn deep semantic patterns from unlabeled data, developing broad language understanding and general-purpose reasoning capabilities.
@@
This study builds on these insights by using the pretrained LLaMA-3.2-3B model as the foundation for a specialized statistical LLM. The model is built on an optimized Transformer decoder-only architecture, featuring several key improvements over the original Transformer design [29]. Whereas the traditional Transformer employs a dual-branch encoder–decoder structure with post-layer normalization, LLaMA retains only the decoder and applies pre-normalization using RMSNorm [68] at the input of each sub-layer, enhancing stability in deep model training. For positional encoding, LLaMA replaces traditional additive embeddings with rotary position embeddings [59]. The self-attention mechanism in LLaMA employs masked attention and supports both multi-head attention and grouped-query attention [1], depending on the model scale. LLaMA’s feed-forward network adopts the SwiGLU [56] activation function. Collectively, these architectural refinements underpin the high efficiency and performance characteristic of the LLaMA family.
@@ 2.2 Supervised fine-tuning
SFT adapts a pretrained language model to a specific downstream task by learning from labeled data $\mathcal{D}=\{(x,y)\}$, where $x$ is the input and $y$ is the target output. The model parameters $\theta$ are optimized by minimizing the negative log-likelihood (NLL) loss:
@@
While SFT is often implemented as FFT, updating all model parameters, this approach becomes prohibitively expensive for large-scale LLMs and is susceptible to catastrophic forgetting [33, 44, 27]. Earlier work ULMFiT [24] introduced techniques to overcome FFT’s potential risks. ULMFiT’s key strategies included: discriminative fine-tuning (applying different learning rates to different layers, preserving low-level general features), gradual unfreezing (unfreezing layers progressively from top to bottom, allowing adaptation of high-level features before modifying lower-level ones), and slanted triangular learning rates (rapidly increasing the learning rate early in training, then gradually decreasing it to balance adaptation speed with knowledge retention). These methods reduced knowledge degradation and inspired more refined fine-tuning strategies in subsequent large-model training.
@@ 2.3 Parameter-efficient fine-tuning
The core idea of PEFT is to freeze most of the parameters of a pretrained model $\mathbf{\Phi}_{0}$ and introduce only a small set of additional learnable parameters $\mathbf{\Theta}$, typically accounting for less than 1% of the total. During fine-tuning, only $\mathbf{\Theta}$ is updated, while $\mathbf{\Phi}_{0}$ remains unchanged. This approach substantially reduces computational and storage costs and mitigates catastrophic forgetting, as the pretrained parameters are preserved, maintaining the general knowledge acquired during pretraining.
Among various PEFT techniques, LoRA [26] has emerged as one of the most widely used due to its simplicity and effectiveness. LoRA is based on the hypothesis that the weight update $\triangle W$ to a pretrained matrix $W_{0}\in\mathbb{R}^{d\times k}$ can be decomposed into the product of two low-rank matrices:
@@
The success of LoRA has led to numerous extensions. QLoRA [14] combines LoRA with 4-bit quantization to reduce memory requirements, enabling fine-tuning of large models on consumer-grade GPUs. DoRA [42] decomposes weights into magnitude and direction, applying low-rank adaptation only to the directional component to potentially improve performance. Further research has examined the impact of LoRA on different layers and modules, such as attention mechanisms and feed-forward networks [39].
@@ 2.4 Continual pretraining
Supervised fine-tuning alone often cannot provide a model with the depth of domain knowledge required to fully capture specialized terminology, linguistic patterns, and background context. To address this limitation, CoP has been proposed as an effective strategy for domain knowledge enhancement. Its central idea is to continue training a general-purpose pretrained model on large amounts of domain-specific text—typically unlabeled—using SSL objectives. By extending the objectives proven effective in initial pretraining, CoP allows the model to efficiently absorb knowledge from the new corpus.
Two common approaches dominate CoP. Masked language modeling (MLM), popularized by BERT, randomly masks tokens in the input and trains the model to predict them based on context. This promotes bidirectional contextual understanding, crucial for mastering precise domain-specific terminology. Causal language modeling (CLM)/next token prediction (NTP), used in autoregressive models such as GPT, predicts the next token given the preceding sequence, enabling the model to learn domain-specific text generation, terminology sequences, and stylistic coherence. Other SSL variants, such as T5’s denoising objective [52] and BART’s text infilling [36], also provide inspiration for CoP design, though MLM and CLM remain the most widely adopted due to their simplicity and effectiveness.
Recent work has focused on improving CoP’s efficiency and stability. For example, the (re)warming strategy leverages better initialization and learning rate schedules to accelerate adaptation and convergence [18]. The lifelong pretraining framework enables continuous learning from evolving domain corpus while mitigating knowledge forgetting, addressing the challenge of rapidly changing knowledge bases [31].
A key advantage of CoP is that it eliminates the need for costly manual annotation. By immersing the model in large volumes of authentic domain-specific text, CoP enables it to acquire vocabulary, grammar, semantic structures, and even implicit reasoning patterns characteristic of the domain. Empirical evidence shows CoP to be particularly effective in terminology-dense fields, in domains with distinct linguistic styles, or when the target corpus diverges substantially from the original pretraining data [19].
@@ 2.5 Instruction tuning
Instruction tuning is a pivotal technique that transforms LLMs from simple “text-completion engines” into versatile AI assistants capable of understanding user intent and following natural language instructions. Its core mechanism is SFT on large-scale datasets of instruction-input-output pairs that span diverse task types and formats. Through this process, models learn not only to interpret instructions but also to generalize effectively to new, unseen tasks.
@@
A major factor behind instruction tuning’s success is the scale, diversity, and quality of training data. Early approaches reformatted existing NLP datasets—such as question answering, summarization, and translation—into instruction-based formats, unifying hundreds of task types under a single paradigm [43]. Later, automated data generation methods further advanced scalability. Self-Instruct [66] leveraged stronger teacher models to generate new task instructions and responses from a small set of seed instructions, while XL-Instruct [30] extended this approach to build large, high-quality corpora through self-guidance. These methods reduce reliance on manual annotation but introduce challenges such as redundancy, lack of diversity, and stylistic uniformity, which necessitate additional control mechanisms. Other research has incorporated real user-model conversation logs into instruction tuning datasets [60]. Although limited in scale and sensitive in terms of privacy and compliance, this type of data reflects authentic usage scenarios and enhances a model’s responsiveness in open-ended dialogue. Its low cost and high application relevance demonstrate the feasibility of instruction tuning under resource-constrained conditions.
In summary, instruction tuning represents a turning point in LLM development, enabling practical human–computer interaction. Its effectiveness derives not only from fine-tuning strategies but also from the diversity and quality of instruction datasets, which may include reformatted NLP tasks, manual annotations, synthetic data generated by teacher models, and real-world interaction data. Together, these sources equip instruction-tuned models with stronger task comprehension and contextual adaptability, laying a robust “base policy” for subsequent alignment techniques such as RLHF.